A pendulum with a bob of mass m is swinging. If the initial height of the bob is increased from h1 to 2(h1), then what is the velocity change?
Answer #1

Using conservation of energy

Velocity at h1

0.5 m v1^2 = mgh1

v1 = sqrt ( 2 mgh1)

If the height increased to twice

v2 = sqrt ( 2 mg h 1 * 2)

v2 = 1.414* v1

I. E velocity increased by factor of sqrt(2)

=========